If I have travelled 20 km an hour for 50 seconds how much metres did I cover?
Answer by TimothyLamb(4379)

s = d / t
d = s * t
---
d = 20 kph * 50 seconds * (1 hour)/(3600 seconds)= 0.28 miles
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php